OpenStudy (anonymous):

find the derivative for 6e^5x + e^-x^2

hartnn (hartnn):

use the chain rule d/dx e^5x = e^5x d/dx (5x) =.... ? d/dx e^-x^2 - e^-x^2 d/dx (-x^2) =... ?